### v2

# Po-Heng Chou1, Da-Chih Lin2, Hung-Yu Wei2, Walid Saad3, and Yu Tsao1

Abstract
In this paper, a measurement-driven framework is proposed for early radio link failure (RLF) prediction in 5G non-standalone (NSA) railway environments. Using 10 Hz metro-train traces with serving and neighbor-cell indicators, we benchmark six models, namely CNN, LSTM, XGBoost, Anomaly Transformer, PatchTST, and TimesNet, under varied observation windows and prediction horizons. When the observation window is three seconds, TimesNet attains the highest F1 score with a three-second prediction horizon, while CNN provides a favorable accuracy-latency tradeoff with a two-second horizon, enabling proactive actions such as redundancy and adaptive handovers. The results indicate that deep temporal models can anticipate reliability degradations several seconds in advance using lightweight features available on commercial devices, offering a practical path to early-warning control in 5G-based railway systems.
I Introduction
The fifth-generation (5G) wireless network has become a key enabler for emerging mission-critical applications that demand highly reliable communications [1]. Among the vertical sectors benefiting from 5G, railway communications represent one of the most reliability-sensitive domains, as wireless links are responsible for delivering train control messages between onboard and ground controllers [2, 3]. Any delay or loss of such control signals can compromise passenger safety and traffic efficiency [2]. Therefore, ensuring stable and predictive reliability in 5G-based railway systems has become a pressing research objective [3].
While 5G NSA provides high data rates, its dual-connectivity architecture, which couples the Long-Term Evolution (LTE) eNodeB and 5G New Radio (NR) gNodeB, introduces additional control-plane complexity. Frequent link switching between the master and secondary nodes may trigger signaling bursts, configuration failures, or transient disconnections during mobility. These structural dynamics make predictive reliability control a central challenge in realizing mission-critical 5G systems.
Despite the rapid deployment of commercial 5G non-standalone (NSA) networks, numerous empirical studies have revealed that their real-world performance often falls short of the theoretical promises of reliability. Measurements across different urban environments show that 5G networks still suffer from non-negligible packet loss and latency fluctuations when compared with mature LTE networks [4, 5]. Such discrepancies are particularly pronounced in high-mobility scenarios, where frequent cell handovers lead to transient disconnections and unstable throughput. Our previous field experiments along the Taipei Metro confirmed this phenomenon: packet losses and excessive latency events predominantly occurred near station areas with intensive handover activities [6]. By analyzing lower-layer signaling messages captured via MobileInsight [7], we identified that handover-related events are the dominant contributors to reliability degradation in 5G NSA networks.
To mitigate these reliability issues, several studies have explored device-side enhancements to improve the stability of 5G connectivity in railway environments. Most of the previous solutions use recovery mechanisms [8, 9] to maintain redundant connections and avoid unnecessary handovers, thereby reducing packet loss and latency. Nevertheless, such methods remain inherently reactive, as they attempt to recover from reliability degradation after it has occurred, without the capability to foresee impending radio link failures (RLFs). In mission-critical railway systems, this lack of early-warning mechanisms poses a persistent risk because communication outages could still occur before mitigation procedures are triggered.
Unlike prior simulation-based or synthetic RLF prediction studies [19, 20, 21], this work presents the first measurement-driven benchmark for early RLF prediction in real 5G NSA metro-train environments. By leveraging field-collected datasets instead of simulated logs, our framework enables quantitative evaluation of learning-based reliability models under authentic mobility, interference, and signaling conditions.
Motivated by these challenges, this paper proposes a measurement-driven, learning-based framework for early failure detection in 5G NSA railway communications. Our objective is to proactively identify potential RLF events before they cause service disruption, enabling preemptive reliability enhancement such as redundancy activation and adaptive handovers.
Based on the practical measurement data collected from metro railway environments, we systematically evaluate the capability of existing learning models to predict reliability degradation in real-world 5G networks. The dataset is sampled at a rate of 10 Hz, corresponding to ten measurement points per second, where each sample includes key radio signal indicators such as the reference signal received power (RSRP) and reference signal received quality (RSRQ) from both serving and neighboring cells. These features capture instantaneous channel conditions as well as short-term fluctuations caused by mobility and handover activities, thereby providing sufficient temporal granularity for modeling reliability dynamics.
In this paper, six representative models are considered, including convolutional neural network (CNN) [10], long short-term memory (LSTM) [11], XGBoost [12], Anomaly Transformer [13], PatchTST [14], and TimesNet [15]. Each model is trained and evaluated under multiple observation windows (1 s, 2 s, 3 s), prediction horizons (1 s, 2 s, 3 s), and sampling schemes with one, two, or three temporal points, either continuous or non-continuous. Rather than proposing a new architecture, we benchmark six representative models under multiple temporal settings, revealing the trade-offs between prediction horizon, observation context, and early-warning reliability in real metro environments.
The main contributions of this paper are summarized as follows:
  • Measurement-driven early warning: We build a supervised framework for proactive RLF prediction using 10 Hz real-world measurements, combining physical-layer indicators (RSRP, RSRQ) with protocol cues to capture reliability dynamics under mobility.
  • Systematic benchmarking: Six representative models (CNN, LSTM, XGBoost, Anomaly Transformer, PatchTST, TimesNet) are evaluated under multiple $(T_{s},T_{p})$ settings. TimesNet yields the best overall F1 at $T_{s}=3$ s and $T_{p}=3$ s, while CNN offers a strong accuracy-latency tradeoff at $T_{p}=2$ s.
  • Deployment insights: RLFs account for most packet losses in metro environments and can be predicted up to two seconds ahead, enabling redundancy activation and adaptive handovers in practical 5G railway systems.
Figure 1: Downlink packet loss distribution by event types and occurrence frequency.
II Related Works
II-A Measurement-based Reliability Studies
Empirical studies have revealed that the reliability and latency performance of commercial 5G NSA networks often diverge from the theoretical promises of reliability. Early measurement campaigns showed that, compared with mature LTE systems, 5G networks still exhibit non-negligible packet losses and unstable latency under mobility [4, 5]. To better understand these issues, we conducted extensive field experiments in the Taipei mass rapid transit (MRT) system to analyze the performance of 5G NSA networks under real train operation in our previous work [6]. Our results demonstrated that most packet loss and excessive latency events occur near metro stations, where dense base-station deployment leads to frequent handovers. Statistical analysis further indicated that up to $96\%$ of downlink packet loss rate (PLR) happens during handover-related intervals, confirming that handover is the dominant cause of reliability degradation. These measurement-driven studies established a solid empirical foundation for understanding the reliability bottlenecks of 5G NSA railway communications, motivating the need for predictive mechanisms capable of anticipating RLFs before they occur.
II-B Device-Side Reliability Enhancement
Beyond measurement analysis, several studies have explored device-level techniques to enhance the reliability or throughput of 5G communications in railway scenarios. Most of the previous researches provide a multi-connectivity mechanism [16, 17]. The mechanism enables the UEs to transmit data with multiple replicated paths, thereby reducing the probability of data loss. Those works showed that the proposed method effectively suppresses packet losses and can even achieve zero-loss operation under optimized configurations. Such device-side approaches provide an immediate means to improve network robustness without modifying base-station infrastructure. However, these solutions cannot handle simultaneous link failures caused by the use of the same or highly correlated transmission channels. To ensure safety in mission-critical systems such as railway control, reliability assurance should evolve toward data-driven paradigms that enable early detection of impending RLFs.
II-C Learning-Based Reliability Prediction and Anomaly Detection
Machine learning (ML) has recently emerged as a powerful paradigm for reliability prediction and anomaly detection in 5G networks. A comprehensive survey in [18] reviewed the evolution of ML methods for fault prediction in telecommunication networks, ranging from traditional neural architectures to recent graph- and transformer-based frameworks. Early attempts, such as the deep neural network (DNN) model proposed in [19], demonstrated that RLFs can be effectively predicted from key performance indicators (KPIs) and control-plane signaling collected in 5G radio access networks (RANs). However, these early DNN-based approaches often lacked the ability to capture complex spatial and temporal dependencies inherent in dense 5G deployments.
To address these limitations, advanced deep learning architectures have been proposed. Transformer-based models have shown superior capability in learning long-term temporal relationships, as demonstrated in [20], where lightweight variants such as Linformer and Performer achieved high detection accuracy with reduced computational cost. Further improvement was achieved by combining graph neural networks (GNNs) with transformers, enabling spatial correlation modeling across multiple base stations [21]. These studies highlight the evolution toward more expressive and scalable frameworks for real-time RLF prediction.
In parallel, ML has also been applied to anomaly detection in broader 5G-IoT environments. For instance, [22] introduced a hybrid CNN-LSTM framework that integrates multi-level temporal features to detect traffic anomalies in distributed IoT infrastructures. Such approaches illustrate the general applicability of data-driven reliability modeling across different layers and verticals of 5G systems.
Despite these advances, most existing methods remain validated under simulated or generic network conditions. Their adaptation to mission-critical railway environments, where safety and latency constraints are far more stringent, has not been systematically explored. The present study aims to bridge this gap by establishing a measurement-driven benchmark for early RLF detection in real 5G NSA metro scenarios, enabling quantitative assessment of existing models under realistic mobility and signal fluctuation patterns.
III Measurements and Observations
III-A Experimental Environment and Tools
This work builds upon the experimental framework established in our previous measurement campaigns [6], employing the same devices, software tools, and test route to ensure data consistency and comparability. All experiments were conducted along the Taipei MRT Brown Line, from Xinhai Station to Taipei Zoo Station, a route that spans urban, semi-open, and underground segments. This path offers diverse radio environments representative of dense metropolitan 5G deployments, making it ideal for investigating reliability variations under real mobility conditions.
In each experiment, the UE was implemented using commercial smartphones connected via universal serial bus (USB) to a laptop, which served as the data client. A remote server located at National Taiwan University acted as the communication endpoint. The downlink (DL) and uplink (UL) data streams were transmitted using the user datagram protocol (UDP), with a fixed payload size of 250 bytes and a transmission rate of 200 kbps. This configuration provided a lightweight yet continuous traffic load suitable for reliability evaluation while minimizing protocol overhead.
A custom Android monitoring application was developed to collect key measurement parameters in real time. The tool sampled radio-layer indicators at 10 Hz, corresponding to ten measurement points per second, including the RSRP and RSRQ of both serving and neighboring LTE and 5G cells. These physical-layer metrics were synchronized with Global Positioning System (GPS) coordinates, enabling spatio-temporal correlation between network performance and geographical location. The resulting dataset thus captured both the instantaneous and short-term variations of wireless link quality along the railway path.
For control-plane observation, we employed MobileInsight [7], an open-source tool that extracts LTE and 5G NR signaling events from modem logs. This allowed us to record detailed radio resource control (RRC) procedures, such as connection setup, reconfiguration, and handover events. The integration of transport-layer packet traces with low-layer signaling messages enabled a comprehensive view of how physical and control-layer behaviors jointly influence communication reliability in high-mobility 5G NSA networks.
III-B Observations
Fig. 1 summarizes the distribution of downlink packet loss across various event categories observed in our experiments. The analysis indicates that two specific events, NAS Recovery (NASR) and Medium Access Control Configuration Failure (MCGF), exhibit the highest packet loss per event. When the overall frequency and impact are combined, the five most influential events are MCGF, Master Node Base Handover (MNBH), Secondary Cell Group Modification (SCGM), End Node Base Handover (ENBH), and NASR, in descending order of contribution.
A critical pattern emerges when correlating these events with radio link status: both MCGF and NASR are classified as RLF events. Together, they account for 58.3% of all downlink packet losses, with MCGF alone responsible for over half of the total (54.6%). This finding confirms that RLFs are the dominant source of reliability degradation in 5G NSA railway communications.
These results suggest that device-side mechanisms can mitigate ongoing failures, yet most losses originate from abrupt and unpredictable RLF events. Identifying such events in their early stages would allow controllers or user equipment to trigger redundancy and adaptive procedures before a full breakdown. Developing learning-based early warning is therefore essential for improving the reliability and resilience of 5G railway communications.
IV RLF Prediction Design
To proactively address radio link failure (RLF) events, we formulate the task as a supervised time-series classification problem. As illustrated in Fig. 2, at each time point $t$ (denoted as Now), the predictor takes a sliding observation window of length $T_{s}$ seconds as input, consisting of historical measurements sampled every $I_{s}$ seconds. The model then outputs the probability that an RLF will occur within the subsequent prediction horizon of $T_{p}$ seconds. For binary classification, a sample is labeled $y=1$ if any RLF occurs within $(t,\,t+T_{p}]$, and $y=0$ otherwise. For the multi-interval variant, a one-hot vector indicates the sub-interval of $(t,\,t+T_{p}]$ in which the next RLF appears. RLF samples form the positive class.
Figure 2: RLF predictor architecture using data from observation window $T_{s}$ to predict events in prediction horizon $T_{p}$.
Feature Selection
To capture the comprehensive characteristics of radio link quality, the RLF predictor employs a diverse set of features. The primary inputs include direct signal strength indicators such as RSRP and RSRQ. In addition, protocol-level information derived from RRC measurements and event reports provides critical context on network state transitions. To address spatial variability in railway environments, the feature set is further expanded to include the RSRP, RSRQ, and cell identifiers of the top-$N$ neighboring cells, enabling the anticipation of handovers and the detection of coverage deficiencies.
Handling Class Imbalance
RLF events are inherently rare in typical operational settings. Our experiments, conducted with a sampling interval of $I_{s}=0.1$ s (10 Hz), show an approximate ratio of one RLF sample per 500 non-RLF samples, resulting in a highly imbalanced dataset that complicates accurate prediction. This fine-grained sampling rate was intentionally chosen because network performance metrics can fluctuate at millisecond timescales; a 0.1-second interval allows us to capture these subtle variations more effectively, thereby improving the model’s ability to identify early precursors of RLF events. To address this challenge, we examined several widely used techniques for dealing with imbalanced data, including synthetic minority over-sampling (SMOTE), random downsampling to achieve a target ratio of 30:1 (non-RLF to RLF), and class-weighted loss functions that emphasize the minority class during training.
V Performance Results
V-A Implementation Environment and Model Setup
All experiments were implemented in Python 3.10 using TensorFlow and XGBoost libraries within Visual Studio Code. Model training and evaluation were executed on a 13th Gen Intel® Core™ i5-13400 CPU (10 cores, 16 threads, base 2.5 GHz, boost 4.6 GHz, 12 MB cache, 65 W TDP) without GPU acceleration. Six representative models, including CNN, LSTM, XGBoost, TimesNet, Anomaly Transformer, and PatchTST, were benchmarked under identical training settings for fair comparison. All models adopted the Adam optimizer with an initial learning rate of $1\times 10^{-3}$, batch size of 128, and early stopping based on the validation Area Under Curve (AUC) (patience =10=10). A learning rate reduction strategy (ReduceLROnPlateau) with a minimum learning rate of $1\times 10^{-5}$ and class weighting for imbalance correction were applied. Training was conducted for up to 60 epochs. Detailed layer structures, preprocessing scripts, and hyperparameter configurations are available in the project repository. [^1] Source code and configuration files are available at: https://github.com/BoneZhou/RLF-Prediction.
V-B Evaluation Setup
Unless otherwise stated, we report accuracy, precision, recall, and F1 at the F1-optimal decision threshold selected on the validation set by sweeping $\tau\in\{0.1,0.2,\ldots,0.9\}$. RLF is treated as the positive class. We consider observation windows $T_{s}\in\{1,2,3\}$ s and prediction horizons $T_{p}\in\{1,2,3\}$ s with a sampling rate of $I_{s}=0.1$ s (10 Hz).
V-C Overall Trends Across Horizons
Tables I summarize results for $T_{s}=3$ s. Moving from $T_{p}=1$ s to $T_{p}=2$ s generally improves F1 for most models, indicating that a slightly longer horizon exposes more pre-failure cues without overly diluting short-term dynamics. While gains from $T_{p}=2$ s to $T_{p}=3$ s are modest for several baselines, TimesNet benefits from the longer horizon and reaches its peak performance at $T_{s}=3$ s and $T_{p}=3$ s. Hence, $T_{p}=2$ s remains a favorable operating point for low-latency prediction (e.g., with CNN), whereas $T_{p}=3$ s can yield the highest overall reliability with TimesNet under sufficient context. Moreover, all evaluated models demonstrated excellent discriminative performance, with AUC values exceeding 0.95, suggesting highly accurate prediction of events.
TABLE I: Performance comparison with observation window $T_{s}=3$ s and sampling rate $I_{s}=0.1$ s (10 Hz).
Model $T_{p}$ Accuracy Precision Recall F1
CNN [10] 1 s 0.9806 0.2705 0.8116 0.4058
2 s 0.9935 0.7456 0.9130 0.8208
3 s 0.9873 0.6968 0.8551 0.7679
LSTM [11] 1 s 0.9782 0.2275 0.6957 0.3429
2 s 0.975 0.3870 0.9058 0.5423
3 s 0.9781 0.5374 0.7633 0.6307
XGBoost [12] 1 s 0.9962 0.7284 0.8551 0.7867
2 s 0.9927 0.7235 0.8913 0.7987
3 s 0.9865 0.6782 0.8551 0.7564
Anomaly Transformer [13] 1 s 0.9811 0.2581 0.6957 0.3765
2 s 0.9800 0.4357 0.7609 0.5541
3 s 0.9737 0.4762 0.7246 0.5747
PatchTST [14] 1 s 0.9825 0.2690 0.6667 0.3833
2 s 0.952 0.2276 0.8116 0.3556
3 s 0.9613 0.3421 0.6280 0.4429
TimesNet [15] 1 s 0.9881 0.3621 0.6087 0.4541
2 s 0.9737 0.3735 0.8986 0.5277
3 s 0.9924 0.8265 0.8744 0.8498
(a) CNN [10] with $T_{p}=2$ s
(b) XGBoost [12] with $T_{p}=2$ s
(c) TimesNet [15] with $T_{p}=3$ s
Figure 3: Prediction hits of the top-3 models based on Tables I. Profiles use $T_{s}=3$ s and $I_{s}=0.1$ s; CNN and XGBoost are shown with $T_{p}=2$ s, and TimesNet with $T_{p}=3$ s. Red markers indicate successful predictions, i.e., time slots where an RLF was predicted to occur within the next $T_{p}$ seconds.
V-D Model-wise Comparison
Across configurations (Table I), TimesNet delivers the overall highest F1 of 0.8498 at $T_{s}=3$ s and $T_{p}=3$ s (Table I), showing its advantage in modeling longer-term temporal dependencies. CNN achieves a comparable F1 of 0.8208 at a shorter horizon ($T_{p}=2$ s), indicating better responsiveness for early-stage prediction with lower latency. Runtime measurements indicate that moving from $T_{p}=1$ s to $T_{p}=3$ s increases per-sample inference time only marginally (on the order of tens of microseconds on our CPU setup), while improving F1 for models that benefit from longer temporal context. XGBoost remains consistently strong (F1 $\approx 0.79$) across all horizons, benefiting from its ensemble stability but limited by its inability to capture high-order temporal correlations. Transformer-based models (Anomaly Transformer and PatchTST) show moderate recall but lower precision, reflecting higher sensitivity to transient signal fluctuations that increase false alarms under short windows. The LSTM baseline performs reasonably well but still lags behind CNN and TimesNet, consistent with the known limitations of recurrent architectures in capturing fast signal variations at 10 Hz sampling rates.
V-E Effect of Observation Window $T_{s}$
Comparing Tables I, enlarging the observation window from $T_{s}=1$ s to $T_{s}=3$ s consistently improves performance for deep temporal models. In particular, TimesNet benefits most from longer contexts, achieving its peak F1 at $T_{s}=3$ s and $T_{p}=3$ s, while CNN reaches near-optimal performance already at $T_{p}=2$ s. XGBoost saturates early, indicating that statistical tree ensembles rely primarily on short-term features. These results confirm that convolutional and 2D temporal models exploit broader time dependencies more effectively than sequential or boosting-based methods.
V-F Computational Cost vs. Context Trade-off
Extending the observation window from $T_{s}=1$ s to $T_{s}=3$ s roughly doubles or triples the theoretical floating point operations per second (FLOPs) for all models. For example, the TimesNet and CNN models respectively grow by about $2.8\times$ and $2.9\times$ in FLOPs. However, the measured inference latency increases by only around $0.2$ ms on our CPU platform, demonstrating that longer temporal contexts can be leveraged with negligible runtime cost and an overall favorable accuracy-latency tradeoff.
V-G Per-horizon Highlights
$T_{p}=1$ s
Short horizons favor recall-oriented models such as CNN and XGBoost, yet their precision remains limited due to scarce pre-failure cues within 1 s.
$T_{p}=2$ s
CNN achieves its best balance between accuracy and latency, reaching F1 = 0.8208 at $T_{s}=3$ s. TimesNet and XGBoost follow closely, showing that moderate prediction horizons ($\approx 2$ s) effectively capture early degradation patterns.
$T_{p}=3$ s
TimesNet achieves the highest overall F1 = 0.8498, benefiting from its multi-period 2D temporal blocks that aggregate longer-term variations. This indicates that under sufficient context length, TimesNet provides the most reliable long-horizon early-warning capability among all tested models.
V-H Operating Points and Thresholding
All reported scores use the F1-optimal threshold $\tau$ chosen on the validation set. In safety-critical operation, $\tau$ can be shifted at deployment to favor recall (earlier alarms) at the expense of precision. We found that model ranking remains stable under reasonable $\tau$ shifts, indicating robust comparative behavior across methods.
V-I Time-domain Early-warning Behavior
Figure 3 visualizes time-indexed prediction hits for the top three models at $T_{s}=3$ s. Red markers cluster before RLF timestamps, confirming that alarms are typically raised within the allowed horizon. Intervals without RLFs remain mostly unflagged, reducing alarm fatigue. Table II further quantifies hit policies relevant to operations: requiring any two points (i.e., two alarms within the horizon) preserves near-perfect coverage for CNN and TimesNet while suppressing sporadic single-point spikes; stricter policies (three consecutive points) reduce false positives but may miss faster-onset events. These trade-offs offer a tunable path from research metrics to actionable policies (e.g., trigger redundancy or selective band locking upon two consecutive positives).
V-J Error Analysis
Qualitative inspection of false negatives reveals two common patterns: (i) rapid-degradation cases where RSRP/RSRQ drop abruptly with minimal prelude, and (ii) interference-driven perturbations that mimic normal variability until the final seconds. False positives frequently coincide with transient neighbor dominance changes that do not culminate in RLF. Incorporating richer neighbor-cell context or selective protocol cues (e.g., imminent reconfiguration signals) can mitigate such errors.
V-K Runtime and Deployability
Short windows ($T_{s}\leq 3$ s at 10 Hz) keep inference budgets low. CNN and TimesNet offer favorable trade-offs between accuracy and latency with regular compute patterns amenable to UEs or edge gateways. In practice, we recommend pairing a recall-biased operating point with a confirmation policy (two consecutive positives) to enable proactive actions while controlling nuisance alarms.
TABLE II: Performance evaluation of Fig. 3.
Index CNN [10] XGBoost [12] TimesNet [15]
Any one point $\frac{23}{23}=100\%$ $\frac{23}{23}=100\%$ $\frac{23}{23}=100\%$
Any two points $\frac{23}{23}=100\%$ $\frac{22}{23}=95.7\%$ $\frac{23}{23}=100\%$
Any three points $\frac{21}{23}=91.3\%$ $\frac{20}{23}=87\%$ $\frac{23}{23}=100\%$
Two consecutive points $\frac{17}{23}=73.9\%$ $\frac{15}{23}=65.2\%$ $\frac{16}{23}=69.6\%$
Three consecutive points $\frac{3}{23}=13\%$ $\frac{3}{23}=13\%$ $\frac{5}{23}=21.7\%$
VI Conclusion
This work introduced a measurement-driven framework for early RLF prediction in 5G NSA railway environments. Using 10 Hz metro-train measurements, six learning models were benchmarked across multiple temporal configurations. TimesNet achieved the highest F1 with a 3 s observation window ($T_{s}=3$ s) and a 3 s prediction horizon ($T_{p}=3$ s), while CNN delivered comparable performance with lower latency at $T_{p}=2$ s. The results show that deep temporal models can anticipate reliability degradation several seconds in advance using lightweight radio indicators, enabling actionable early warning for proactive control in railway communications. Future research will extend the proposed framework to multi-cell and multi-train scenarios, where inter-train interference and network-side coordination are jointly considered. Moreover, self-adaptive learning at the edge and transfer learning across different routes and operators will be investigated to maintain prediction robustness under dynamic radio environments while reducing data labeling costs for large-scale deployment.