### v4

# Po-Heng Chou1,3, Da-Chih Lin2, Hung-Yu Wei2, Walid Saad3, and Yu Tsao1

Abstract
This paper presents a measurement-driven study of early warning for reliability breakdown events in 5G non-standalone (NSA) railway networks. Using 10 Hz metro-train measurement traces with serving- and neighbor-cell indicators, we benchmark six representative learning models, including CNN, LSTM, XGBoost, Anomaly Transformer, PatchTST, and TimesNet, under multiple observation windows and prediction horizons. Rather than proposing a new prediction architecture, this study develops a measurement-driven benchmark to quantify the feasibility and operating trade-offs of seconds-ahead reliability prediction in 5G NSA railway environments. Experimental results show that learning models can anticipate RLF-related reliability breakdown events seconds in advance using lightweight radio features available on commercial devices. The presented benchmark provides insights for sensing-assisted communication control and offers an empirical foundation for integrating sensing and analytics into future mobility control.
I Introduction
The fifth-generation (5G) wireless network has become a key enabler for emerging mission-critical applications that demand highly reliable communications [1]. Among the vertical sectors benefiting from 5G, railway communications represent one of the most reliability-sensitive domains, as wireless links are responsible for delivering train control messages between onboard and ground controllers [2, 3]. Any delay or loss of such control signals can compromise passenger safety and traffic efficiency [2]. Therefore, ensuring stable and predictive reliability in 5G-based railway systems has become a pressing research objective [3].
While 5G NSA provides high data rates, its dual-connectivity architecture, which couples the Long-Term Evolution (LTE) eNodeB and 5G New Radio (NR) gNodeB, introduces additional control-plane complexity. Frequent link switching between the master and secondary nodes may trigger signaling bursts, configuration failures, or transient disconnections during mobility. These structural dynamics make predictive reliability control a central challenge in realizing mission-critical 5G systems.
Despite the rapid deployment of commercial 5G non-standalone (NSA) networks, numerous empirical studies have revealed that their real-world performance often falls short of the theoretical promises of reliability. Measurements across different urban environments show that 5G networks still suffer from non-negligible packet loss and latency fluctuations when compared with mature LTE networks [4]. Such discrepancies are particularly pronounced in high-mobility scenarios, where frequent cell handovers lead to transient disconnections and unstable throughput. Our previous field experiments along the Taipei Metro confirmed this phenomenon: packet losses and excessive latency events predominantly occurred near station areas with intensive handover activities [5]. By analyzing lower-layer signaling messages captured via MobileInsight [6], we identified that handover-related events are the dominant contributors to reliability degradation in 5G NSA networks. From a 6G viewpoint, early RLF warning can be regarded as device-side network sensing that supports sensing-assisted mobility control. In this context, mobile devices can be viewed as distributed sensing nodes that continuously monitor network reliability under high-mobility conditions.
To mitigate these reliability issues, several studies have explored device-side enhancements to improve the stability of 5G connectivity in railway environments. Most prior solutions rely on recovery mechanisms [7, 8] to maintain redundant connections and avoid unnecessary handovers, thereby reducing packet loss and latency. While such operational studies provide important insights into failure handling and recovery in commercial 5G deployments, they remain fundamentally reactive because mitigation is triggered only after degradation has emerged. In contrast, this work focuses on the complementary problem of seconds-ahead early warning using lightweight radio features on commercial devices. The goal of this paper is therefore not to replace failure recovery mechanisms, but to quantify the feasibility and operating trade-offs of predictive warning before reliability breakdown occurs.
Unlike prior simulation-based or synthetic RLF prediction studies [18, 19, 20], this work presents a measurement-driven benchmarking study for early prediction of RLF-related reliability breakdown events in a real 5G NSA metro-train environment. By leveraging field-collected datasets instead of simulated logs, the proposed framework enables a quantitative evaluation of learning-based reliability models under authentic mobility, interference, and signaling conditions.
Motivated by these challenges, this paper presents a measurement-driven framework for early detection of reliability breakdown events in 5G NSA railway communications. Our objective is to identify impending reliability breakdown events before they cause observable service disruption, thereby supporting preemptive reliability enhancement such as warning generation, redundancy preparation, and adaptive handover assistance.
Based on the practical measurement data collected from metro railway environments, we systematically evaluate the capability of existing learning models to predict reliability degradation in real-world 5G networks. The dataset is sampled at a rate of 10 Hz, corresponding to ten measurement points per second, where each sample includes key radio signal indicators such as the reference signal received power (RSRP) and reference signal received quality (RSRQ) from both serving and neighboring cells. These features capture instantaneous channel conditions as well as short-term fluctuations caused by mobility and handover activities, thereby providing sufficient temporal granularity for modeling reliability dynamics. We benchmark six representative models, namely CNN [9], LSTM [10], XGBoost [11], Anomaly Transformer [12], PatchTST [13], and TimesNet [14], under multiple observation windows and prediction horizons, where $T_{s}$ and $T_{p}$ denote the observation window and prediction horizon, respectively.
The key contributions of this study are summarized as follows:
  • Measurement-driven early-warning benchmark based on device-side sensing: We present a supervised early-warning framework using 10 Hz real-world railway measurements collected from a commercial 5G NSA metro environment, leveraging lightweight radio indicators (RSRP, RSRQ) and neighbor-cell dynamics.
  • Systematic model comparison under temporal design choices: Six representative models (CNN, LSTM, XGBoost, Anomaly Transformer, PatchTST, and TimesNet) are evaluated under multiple $(T_{s},T_{p})$ settings to characterize the trade-offs among sensing context, prediction horizon, and alarm reliability.
  • Operational insights for proactive mobility control: We show that RLF-related reliability breakdown events can be anticipated seconds ahead, and we analyze deployment-relevant trigger policies that can support proactive actions such as redundancy activation and adaptive handovers.
Figure 1: Downlink packet loss distribution by event types and occurrence frequency.
II Related Works
II-A Measurement-based Reliability Studies
Empirical studies have revealed that the reliability and latency performance of commercial 5G NSA networks often diverge from the theoretical promises of reliability. Early measurement campaigns showed that, compared with mature LTE systems, 5G networks still exhibit non-negligible packet losses and unstable latency under mobility [4]. To better understand these issues, we conducted extensive field experiments in the Taipei mass rapid transit (MRT) system to analyze the performance of 5G NSA networks under real train operation in our previous work [5]. Our results demonstrated that most packet loss and excessive latency events occur near metro stations, where dense base-station deployment leads to frequent handovers. Statistical analysis further indicated that up to $96\%$ of downlink packet loss rate happens during handover-related intervals, confirming that handover is the dominant cause of reliability degradation. These measurement-driven studies established a solid empirical foundation for understanding the reliability bottlenecks of 5G NSA railway communications, motivating the need for predictive mechanisms capable of anticipating RLFs before they occur.
II-B Device-Side Reliability Enhancement
Beyond measurement analysis, several studies have explored device-level techniques to enhance the reliability or throughput of 5G communications in railway scenarios. Representative approaches include multi-connectivity and recovery-oriented mechanisms [15, 16, 7, 8], which maintain redundant connections, avoid unnecessary handovers, or recover from degraded links after failure symptoms emerge. These methods provide an immediate means to improve network robustness without modifying base-station infrastructure. However, they are still fundamentally reactive and may not prevent abrupt reliability breakdown before mitigation is triggered. This motivates data-driven early-warning mechanisms that can complement recovery-oriented designs by predicting impending failures in advance.
II-C Learning-Based Reliability Prediction and Anomaly Detection
Machine learning (ML) has recently emerged as a powerful paradigm for reliability prediction and anomaly detection in 5G networks. A comprehensive survey in [17] reviewed the evolution of ML methods for fault prediction in telecommunication networks, ranging from traditional neural architectures to recent graph- and transformer-based frameworks. Early attempts, such as the deep neural network (DNN) model proposed in [18], demonstrated that RLFs can be effectively predicted from key performance indicators (KPIs) and control-plane signaling collected in 5G radio access networks (RANs). However, these early DNN-based approaches often lacked the ability to capture complex spatial and temporal dependencies inherent in dense 5G deployments.
To address these limitations, advanced deep learning architectures have been proposed. Transformer-based models have shown superior capability in learning long-term temporal relationships, as demonstrated in [19], where lightweight variants such as Linformer and Performer achieved high detection accuracy with reduced computational cost. Further improvement was achieved by combining graph neural networks (GNNs) with transformers, enabling spatial correlation modeling across multiple base stations [20]. These studies highlight the evolution toward more expressive and scalable frameworks for real-time RLF prediction.
In parallel, ML has also been applied to anomaly detection in broader 5G-IoT environments. For instance, [21] introduced a hybrid CNN-LSTM framework that integrates multi-level temporal features to detect traffic anomalies in distributed IoT infrastructures. Such approaches illustrate the general applicability of data-driven reliability modeling across different layers and verticals of 5G systems. Overall, these studies motivate a measurement-grounded view of early RLF warning as network sensing and analytics that can support sensing-assisted mobility control in future integrated communication systems.
III Measurements and Observations
III-A Experimental Environment and Tools
This work builds upon the experimental framework established in our previous measurement campaigns [5], employing the same devices, software tools, and test route to ensure data consistency and comparability. All experiments were conducted along the Taipei MRT Brown Line, from Xinhai Station to Taipei Zoo Station, a route that spans urban, semi-open, and underground segments. This path offers diverse radio environments representative of dense metropolitan 5G deployments, making it ideal for investigating reliability variations under real mobility conditions.
In each experiment, the UE was implemented using commercial smartphones connected via universal serial bus (USB) to a laptop, which served as the data client. A remote server located at National Taiwan University acted as the communication endpoint. The downlink and uplink data streams were transmitted using the user datagram protocol (UDP), with a fixed payload size of 250 bytes and a transmission rate of 200 kbps. This configuration provided a lightweight yet continuous traffic load suitable for reliability evaluation while minimizing protocol overhead.
A custom Android monitoring application was developed to collect key measurement parameters in real time. The tool sampled radio-layer indicators at 10 Hz, corresponding to ten measurement points per second, including the RSRP and RSRQ of both serving and neighboring LTE and 5G cells. These physical-layer metrics were synchronized with Global Positioning System (GPS) coordinates, enabling spatio-temporal correlation between network performance and geographical location. The resulting dataset thus captured both the instantaneous and short-term variations of wireless link quality along the railway path.
For control-plane observation, we employed MobileInsight [6], an open-source tool that extracts LTE and 5G NR signaling events from modem logs. This allowed us to record detailed radio resource control (RRC) procedures, such as connection setup, reconfiguration, and handover events. The integration of transport-layer packet traces with low-layer signaling messages enabled a comprehensive view of how physical and control-layer behaviors jointly influence communication reliability in high-mobility 5G NSA networks. Although the campaign was conducted on a single metro line and operator setup, the route spans urban, semi-open, and underground segments, providing heterogeneous propagation conditions for an initial measurement-driven benchmark.
III-B Observations
Fig. 1 summarizes the distribution of downlink packet loss across various event categories observed in our experiments. The analysis indicates that two specific events, Non-Access Stratum Recovery (NASR) and Master Cell Group Failure (MCGF), exhibit the highest packet loss per event. When the overall frequency and impact are combined, the five most influential events are MCGF, Master eNB Handover (MNBH), Secondary Cell Group Modification (SCGM), eNB Handover (ENBH), and NASR, in descending order of contribution.
A critical pattern emerges when correlating these events with radio link reliability: both MCGF and NASR correspond to reliability breakdown events that are strongly associated with radio link failure behavior in the measured 5G NSA network. Together, they account for 58.3% of all downlink packet losses, with MCGF alone responsible for over half of the total (54.6%). This finding confirms that RLF-related reliability breakdown events are the dominant source of reliability degradation in the measured 5G NSA railway environment.
These observations indicate that many reliability breakdown events are preceded by measurable radio-layer fluctuations, suggesting that lightweight device-side indicators can support early warning before a full breakdown. Developing learning-based early warning is therefore essential for improving the reliability and resilience of 5G railway communications.
IV RLF Prediction Design
In this paper, the positive class is defined using stack-level events observed from operational logs, primarily MCGF and NASR. These events do not strictly correspond to the formal 3GPP RLF definition based on out-of-sync indications and timer expiry. However, they represent practically observable reliability breakdown events that lead to packet loss or service interruption in commercial 5G NSA operation. Therefore, the prediction target in this work is more precisely interpreted as RLF-related reliability breakdown events. To proactively address RLF events, we formulate the task as a supervised time-series classification problem. As illustrated in Fig. 2, at each time point $t$ (denoted as Now), the predictor takes a sliding observation window of length $T_{s}$ seconds as input, consisting of historical measurements sampled every $I_{s}$ seconds. The model then outputs the probability that an RLF-related reliability breakdown event will occur within the subsequent prediction horizon of $T_{p}$ seconds. For binary classification, a sample is labeled $y=1$ if any target reliability breakdown event occurs within $(t,\,t+T_{p}]$, and $y=0$ otherwise. For the multi-interval variant, a one-hot vector indicates the sub-interval of $(t,\,t+T_{p}]$ in which the next RLF appears. Positive samples correspond to RLF-related reliability breakdown events.
Figure 2: RLF predictor architecture using data from observation window $T_{s}$ to predict events in prediction horizon $T_{p}$.
Feature Selection
To capture the comprehensive characteristics of radio link quality, the RLF predictor employs a diverse set of features. The primary inputs include direct signal strength indicators such as RSRP and RSRQ. In addition, protocol-level information derived from RRC measurements and event reports provides critical context on network state transitions. To address spatial variability in railway environments, the feature set is further expanded to include the RSRP, RSRQ, and cell identifiers of the top-$N$ neighboring cells, enabling the anticipation of handovers and the detection of coverage deficiencies.
Handling Class Imbalance
RLF events are inherently rare in typical operational settings. Our experiments, conducted with a sampling interval of $I_{s}=0.1$ s (10 Hz), show an approximate ratio of one RLF sample per 500 non-RLF samples, resulting in a highly imbalanced dataset that complicates accurate prediction. This fine-grained sampling rate was intentionally chosen because network performance metrics can fluctuate at millisecond timescales; a 0.1-second interval allows us to capture these subtle variations more effectively, thereby improving the model’s ability to identify early precursors of RLF events. To mitigate this challenge, we considered several widely used strategies for imbalanced learning, including synthetic minority over-sampling (SMOTE), random downsampling to achieve a target ratio of 30:1 (non-RLF to RLF), and class-weighted loss functions that emphasize the minority class during training.
V Performance Results
(a) CNN [9] with $T_{p}=2$ s
(b) XGBoost [11] with $T_{p}=2$ s
(c) TimesNet [14] with $T_{p}=3$ s
Figure 3: Prediction hits of the top-3 models based on Tables I. Profiles use $T_{s}=3$ s and $I_{s}=0.1$ s; CNN and XGBoost are shown with $T_{p}=2$ s, and TimesNet with $T_{p}=3$ s. Red markers indicate correct predictions, i.e., time slots where an RLF-related reliability breakdown event was predicted to occur within the next $T_{p}$ seconds.
V-A Implementation Environment and Model Setup
All experiments were implemented in Python 3.10 using TensorFlow and XGBoost libraries within Visual Studio Code. Model training and evaluation were executed on a 13th Gen Intel® Core™ i5-13400 CPU (10 cores, 16 threads, base 2.5 GHz, boost 4.6 GHz, 12 MB cache, 65 W TDP) without GPU acceleration. Six representative models, including CNN, LSTM, XGBoost, TimesNet, Anomaly Transformer, and PatchTST, were benchmarked under identical training settings for fair comparison. All models adopted the Adam optimizer with an initial learning rate of $1\times 10^{-3}$, batch size of 128, and early stopping based on the validation Area Under Curve (AUC) (patience =10=10). A learning rate reduction strategy (ReduceLROnPlateau) with a minimum learning rate of $1\times 10^{-5}$ and class weighting for imbalance correction were applied. Training was conducted for up to 60 epochs. Detailed layer structures, preprocessing scripts, and hyperparameter configurations are available in the project repository. [^1] Source code and configuration files are available at: github.com/BoneZhou/RLF-Prediction.
V-B Evaluation Setup
Unless otherwise stated, we report accuracy, precision, recall, and F1 at the F1-optimal decision threshold selected on the validation set by sweeping $\tau\in\{0.1,0.2,\ldots,0.9\}$. The positive class corresponds to RLF-related events. We consider observation windows $T_{s}\in\{1,2,3\}$ s and prediction horizons $T_{p}\in\{1,2,3\}$ s with a sampling rate of $I_{s}=0.1$ s (10 Hz).
Because the dataset is highly imbalanced (approximately 1:500 positive-to-negative ratio), accuracy and ROC-oriented metrics alone may overestimate practical utility. Therefore, in addition to accuracy, precision, recall, and F1, we also evaluate deployment-oriented alarm reliability using PR-AUC, false alarms per minute, and recall under fixed false-alarm budgets whenever applicable. These metrics better reflect the operational usefulness of early-warning decisions in highly imbalanced railway communication settings.
Beyond sample-level classification scores, an early-warning system should also be assessed from an operational perspective. In particular, excessive false alarms may trigger unnecessary redundancy activation, handover attempts, or control signaling overhead in safety-critical railway systems. Therefore, we additionally consider alarm-centric metrics such as PR-AUC, false alarms per minute, and event-level hit rate under different trigger policies.
V-C Overall Trends Across Horizons
Tables I summarize results for $T_{s}=3$ s. Moving from $T_{p}=1$ s to $T_{p}=2$ s generally improves F1 for most models, indicating that a slightly longer horizon exposes more pre-failure cues without overly diluting short-term dynamics. While gains from $T_{p}=2$ s to $T_{p}=3$ s are modest for several baselines, TimesNet benefits from the longer horizon and reaches its peak performance at $T_{s}=3$ s and $T_{p}=3$ s. Hence, $T_{p}=2$ s remains a favorable operating point for low-latency prediction (e.g., with CNN), whereas $T_{p}=3$ s can yield the highest overall reliability with TimesNet under sufficient context. Moreover, all evaluated models achieved AUC values exceeding 0.95, indicating strong ranking capability, although such ROC-oriented metrics should be interpreted together with PR-AUC and alarm-centric measures under severe class imbalance.
V-D Model-wise Comparison
Across configurations (Table I), TimesNet delivers the overall highest F1 of 0.8498 at $T_{s}=3$ s and $T_{p}=3$ s (Table I), showing its advantage in modeling longer-term temporal dependencies. CNN achieves a comparable F1 of 0.8208 at a shorter horizon ($T_{p}=2$ s), indicating better responsiveness for early-stage prediction with lower latency. Runtime measurements indicate that moving from $T_{p}=1$ s to $T_{p}=3$ s increases per-sample inference time only marginally (on the order of tens of microseconds on our CPU setup), while improving F1 for models that benefit from longer temporal context. XGBoost remains consistently strong (F1 $\approx 0.79$) across all horizons, benefiting from its ensemble stability but limited by its inability to capture high-order temporal correlations. Transformer-based models (Anomaly Transformer and PatchTST) show moderate recall but lower precision, reflecting higher sensitivity to transient signal fluctuations that increase false alarms under short windows. The LSTM baseline performs reasonably well but still lags behind CNN and TimesNet, consistent with the known limitations of recurrent architectures in capturing fast signal variations at 10 Hz sampling rates.
V-E Effect of Observation Window $T_{s}$
Comparing Tables I, enlarging the observation window from $T_{s}=1$ s to $T_{s}=3$ s consistently improves performance for deep temporal models. In particular, TimesNet benefits most from longer contexts, achieving its peak F1 at $T_{s}=3$ s and $T_{p}=3$ s, while CNN reaches near-optimal performance already at $T_{p}=2$ s. XGBoost saturates early, indicating that statistical tree ensembles rely primarily on short-term features. These results confirm that convolutional and 2D temporal models exploit broader time dependencies more effectively than sequential or boosting-based methods.
TABLE I: Performance comparison with observation window $T_{s}=3$ s and sampling rate $I_{s}=0.1$ s (10 Hz).
Model $T_{p}$ Accuracy Precision Recall F1
CNN [9] 1 s 0.9806 0.2705 0.8116 0.4058
2 s 0.9935 0.7456 0.9130 0.8208
3 s 0.9873 0.6968 0.8551 0.7679
LSTM [10] 1 s 0.9782 0.2275 0.6957 0.3429
2 s 0.975 0.3870 0.9058 0.5423
3 s 0.9781 0.5374 0.7633 0.6307
XGBoost [11] 1 s 0.9962 0.7284 0.8551 0.7867
2 s 0.9927 0.7235 0.8913 0.7987
3 s 0.9865 0.6782 0.8551 0.7564
Anomaly Transformer [12] 1 s 0.9811 0.2581 0.6957 0.3765
2 s 0.9800 0.4357 0.7609 0.5541
3 s 0.9737 0.4762 0.7246 0.5747
PatchTST [13] 1 s 0.9825 0.2690 0.6667 0.3833
2 s 0.952 0.2276 0.8116 0.3556
3 s 0.9613 0.3421 0.6280 0.4429
TimesNet [14] 1 s 0.9881 0.3621 0.6087 0.4541
2 s 0.9737 0.3735 0.8986 0.5277
3 s 0.9924 0.8265 0.8744 0.8498
V-F Computational Cost vs. Context Trade-off
Extending the observation window from $T_{s}=1$ s to $T_{s}=3$ s roughly doubles or triples the theoretical FLOPs for all models. However, the measured inference latency increases by only about $0.2$ ms on our CPU platform, indicating that longer temporal contexts can be exploited with negligible runtime penalty.
V-G Time-domain Early-warning Behavior
Figure 3 visualizes time-indexed prediction hits for the top three models at $T_{s}=3$ s. Red markers cluster before labeled reliability-breakdown timestamps, confirming that alarms are typically raised within the allowed horizon. Intervals without labeled breakdown events remain mostly unflagged, suggesting that nuisance alarms can be partially controlled under suitable trigger policies. Table II further quantifies hit policies relevant to operations: requiring any two points (i.e., two alarms within the horizon) preserves near-perfect coverage for CNN and TimesNet while suppressing sporadic single-point spikes; stricter policies (three consecutive points) reduce false positives but may miss faster-onset events. These trade-offs provide a tunable path from research metrics to deployment-oriented alarm policies, for example, issuing a proactive warning only after two consecutive positive decisions.
TABLE II: Performance evaluation of Fig. 3.
Index CNN [9] XGBoost [11] TimesNet [14]
Any one point $\frac{23}{23}=100\%$ $\frac{23}{23}=100\%$ $\frac{23}{23}=100\%$
Any two points $\frac{23}{23}=100\%$ $\frac{22}{23}=95.7\%$ $\frac{23}{23}=100\%$
Any three points $\frac{21}{23}=91.3\%$ $\frac{20}{23}=87\%$ $\frac{23}{23}=100\%$
Two consecutive points $\frac{17}{23}=73.9\%$ $\frac{15}{23}=65.2\%$ $\frac{16}{23}=69.6\%$
Three consecutive points $\frac{3}{23}=13\%$ $\frac{3}{23}=13\%$ $\frac{5}{23}=21.7\%$
V-H Error Analysis
Qualitative inspection shows that false negatives mainly arise from abrupt degradation with little prelude or interference-driven perturbations that resemble normal variability, while false positives often coincide with transient neighbor-dominance changes. These observations motivate confirmation-based trigger policies and false-alarm-budget evaluation in practical railway settings.
VI Conclusion
This work presented a measurement-driven benchmark for early prediction of RLF-related reliability breakdown events in 5G NSA railway environments. Using 10 Hz metro-train measurements, six learning models were evaluated across multiple temporal configurations. TimesNet achieved the best overall F1 at $(T_{s},T_{p})=(3\,\mathrm{s},3\,\mathrm{s})$, while CNN offered a strong latency-reliability trade-off at $T_{p}=2$ s. The results suggest that lightweight device-side radio indicators can support seconds-ahead reliability warning for IoT-enabled transportation systems. Future work will extend the benchmark to multi-line and cross-environment settings and further investigate event-level and false-alarm-budget evaluation.